OpenStudy (anonymous):

write the equation of the locus of a point, the square of whose distance from (-2,-5) is three times its distance from the line 8x+15y=34

OpenStudy (anonymous):



OpenStudy (campbell_st):

well you have a point P(x, y) and its distance from (-2, -5) is \[d = \sqrt{x +2)^2 + (y + 5)^2}\] now square it \[d^2 = (x + 2)^2 + (y + 5)^2\] does that make sense...?

OpenStudy (anonymous):

x^2 +4x + y^2 +10y +29 = ....

OpenStudy (campbell_st):

now you need the perpendiculat distance from a point to a line... point P(x, y) and the line 8x + 15y - 34 = 0 so \[l = \frac{\left| 8x + 15y - 34 \right|}{\sqrt{8^2 + 15^2}}\] then it says d^2 = 3l so equate them and then keep working on it... hope it makes sense